OpenStudy (anonymous):

sum n=0 to inf [(2n)!/(n)!^2]*(1/4)^n

OpenStudy (anonymous):

\[\sum_{n=0}^{+\infty}\bigg(\frac{(2n)!}{n!^2}\bigg)\bigg(\frac{1}{4}\bigg)^n\] I don't have the answer but this could help people who can actually answer it...

OpenStudy (anonymous):

i hope so...all the divergence tests i know failed :(

OpenStudy (anonymous):

i need to proof that the series diverges